OpenStudy (math2400):

Calc prob. I know how to solve using d(t) = vt + 16t^2 but that's not the point of the assignment. I need to somehow incorporate derivatives/integrals so could someone explain how I would do that please?

OpenStudy (math2400):

An object is thrown off of a 480ft building with initial velocity of -64 ft/sec. With what velocity does the object hit the ground?

OpenStudy (anonymous):

starting with gravity?

OpenStudy (anonymous):

gravity is a constant, \(-32\) feat per second squared i.e. it is acceleration it is the derivative of the velocity is \(-32\) so \[v=32t+ v_0\] as you know \(v(0)=-64\) you have \(v_0=-64\) and so \[v(t)=-32t-64\]

OpenStudy (anonymous):

feet, not feat

OpenStudy (anonymous):

then the position is the anti derivative of the velocity, so \[h(t)=-16t^2-64t+h_0\] and since \(h(0)=480\) you know \[h(t)=-16t^2-64t+480\]

OpenStudy (anonymous):

it hits the ground when \(h(t)=0\) solve \(-16t^2-64t+480=0\) and get \(t=-2+\sqrt{34}\)

OpenStudy (anonymous):

finally to find the velocity when it hits the ground, compute \[v(2-\sqrt{34})\]

OpenStudy (math2400):

Thank yoU!!! I completely forgot about this! You're the best!